A block slides down a smooth ramp, starting from rest at a height h. When it reaches the bottom it’s moving at speed v. It then continues to slide up a second smoothramp.

At what height is its speed equal to v/2?

a.)h/4
b.)h/2
c.)3h/4
d.)2h

Please show which formula you used.
Answer #1

ΔE = 0


m g h - ½ m vi² = 0


vi² = 2 g h


and then, continue slide up to the second ramp,


ΔE = 0


-m g h' - ½ m v² +  ½ m vi² = 0


v = ½ vi and canceled out of mass "m"


-g h' - ½ (½ vi)² +  ½ vi² = 0


(3/8) vi² = g h'


but please remember that :


2 g h = vi²


(3/8) (2 g h) = g h'


h' = (3/4) h


source: Mastering Physics
answered by: Aziz